Myers Business Systems is evaluating the introduction of a new product. The possible levels of unit sales and the probabilities of their occurrence are given next:

Possible Market Reaction Sales in Units Probabilities
Low response 25 0.20
Moderate response 40 0.30
High response 65 0.20
Very high response 80 0.30

a. What is the expected value of unit sales for the new product? (Do not round intermediate calculations and round your answer to the nearest whole unit.)

b. What is the standard deviation of unit sales? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Expected value=Respective sales*Respective probability

=(0.2*25)+(0.3*40)+(0.2*65)+(0.3*80)=54

probability Sales probability*(Sales-Expected value)^2
0.2 25 0.2*(25-54)^2=168.2
0.3 40 0.3*(40-54)^2=58.8
0.2 65 0.2*(65-54)^2=24.2
0.3 80 0.3*(80-54)^2=202.8
Total=454

Standard deviation=[Total probability*(Sales-Expected value)^2/Total probability]^(1/2)

=(454)^(1/2)

=21.31(Approx)